OpenStudy (anonymous):

you parents have $20,000 in the bank for your college education. in how many years can they have the amount of $60,000 to pay for college expends if the annual interest is 6% and it is compounded continuously? *formula: Q(t)= Q initial e^rt

OpenStudy (anonymous):

Q(t)=60,000 Q= 20,000 r=0.06 Divide both sides by 20,000... Then you can ln both sides to get rid of the e.

OpenStudy (anonymous):

can you show me how to do that please?

OpenStudy (anonymous):

whats the eqation?

OpenStudy (anonymous):

60,000 = 20,000e^(0.06t) 3=e^(0.06t) ln 3=lne^(0.06t)

OpenStudy (anonymous):

\[60000 = 20000 e ^{0.06t}\] \[60000/20000 = 20000/20000 e ^{0.06t}\] \[3 = e ^{0.06t}\] \[ln(3) = 0.06t}\] \[ln(3)/0.06 = t}\]

OpenStudy (anonymous):

so the final answer is ln(3)/0.06 = t}?

OpenStudy (anonymous):

\[t = ln(3)/0.06\] yes

OpenStudy (anonymous):

my calculator says 18.3

OpenStudy (anonymous):

i have a TI-84 graphong calculator and my teacher didnt rlly teach me how to use it. so wheres the ln button?

OpenStudy (anonymous):

never mind got it!

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

I will need one of those graphing calculators for my next sem

OpenStudy (anonymous):

what math are you taking atm?

OpenStudy (anonymous):

advance calculus and differential equations, starting sept.

OpenStudy (anonymous):

sounds scary... >.<